OpenStudy (anonymous):

Malorie can work 10 math problems in 3 minutes. At this rate, how many problems can she work in one hour?

OpenStudy (anonymous):

Would it be 200?

OpenStudy (acxbox22):

same thing...use proportions

OpenStudy (acxbox22):

10/3=x/60 3x=600 x=200 you are right

OpenStudy (anonymous):

oh yay

OpenStudy (whpalmer4):

Find the rate. X problems/y time is x/y problems/time. Multiply by time to get problems worked. For example, 3 problems in 5 minutes is a rate of 3/5 problems/minute. In 25 minutes, 3/5 problems/minute*25 minutes = 75/5 problems = 15 problems. I find computing the rate safer (though equivalent) than doing proportions, especially if unit conversions are required.

OpenStudy (yolomcswagginsggg):

;-;